# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

### v4

@@ 2.4 Numerical Simulations
1 (4,1) $(97.4\%,\ 5.3\times 10^{-2})$ (6,1) $(98.4\%,\ 7.5\times 10^{-2})$ (8,1) $(98.6\%,\ 9.5\times 10^{-2})$
@@
Table 3: Performance of an ESPD initialized with a SPD from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
@@
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
(d) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
Figure 2: Performance of ESPD initialized with SPDs from [28] ((a), (b)) and [29] ((c), (d)), respectively.
2.5 Implications for QKD
As a representative application scenario, the implications of the ESPD framework for QKD are analyzed. For a QKD protocol characterized by a secure error threshold $e_{th}$ and a total error rate $e_{C}$ excluding detector contributions, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by[27]:
Since the ESPD paradigm can substantially suppress the effective DCR $d$, the minimal tolerable channel transmission rate $\gamma$ is proportionately reduced. Furthermore, the simultaneous enhancement of the DE $\eta$ not only reduces $\gamma$ but also directly improves the achievable key rate at a fixed channel transmission rate.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD framework achieves this improvement without introducing additional sifting procedures that reduce the effective sifting efficiency. Here, QKD performance can be enhanced through the simultaneous suppression of the effective DCR and enhancement of the effective DE, without the need for auxiliary sifting operations. In this sense, the ESPD paradigm provides a more efficient system-level approach for improving QKD performance under practical constraints.
3 Discussion
In the absence of an experimental demonstration, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations as well as directions for future work.
3.1 Technological Feasibility and Performance under Degraded Parameters
The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs. The device requirements are discussed as follows.
3.1.1 State Preparations and Measurements
The flexibility of the ESPD architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be achieved at extremely low levels. Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
@@ 3.1.2 Controlled Gates
Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96. Simulations are performed for $(n,k)=(5,2)$ at each level, and all other settings remain consistent with those described in Section 2.4.
@@
Figure 3: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, does not decrease with an increasing number of cascaded modules or ESPD levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
@@
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.3 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
@@
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryonic components[39, 40, 41, 42]. However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), with visual representation in Fig.4. All other settings remain consistent with those described in Section 2.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with existing room-temperature quantum technologies.
@@
Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
@@
Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
@@
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.40$.
(d) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$, with $P=0.40$.
Figure 4: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$ ((a), (b)) and $P=0.40$ ((c), (d)), respectively.
3.1.3 Choice of initialized SPDs
Notably, the requirements imposed on the SPD used for initialization are minimal, which further lowers the practical barrier to implementation. As evidenced by the comparison between Tables 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the initialized SPD, provided that it lies within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, numerical simulations suggest that both SPDs with low DE and low DCR[29], as well as those with moderate DE and relatively high DCR[28], can serve as viable initializations.
In summary, the ESPD scheme requires only a conventional room-temperature SPD for initialization, which is routinely achievable with existing technology. For instance, the SPD developed two decades ago [28] is sufficient.
@@ 3.1.4 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
3.2 Limitations and Future Works
@@ 3.2.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
3.2.2 Resource Overhead vs. Environmental Simplicity
@@
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
@@
However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
@@
In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation. The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K). The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration. By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost. This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.

### v5

@@ 2.4 Monte Carlo Steady-State Analysis and Numerical Simulations
1 (4,2) $(97.4\%,\ 5.3\times 10^{-2})$ (6,3) $(98.4\%,\ 7.5\times 10^{-2})$ (8,4) $(98.6\%,\ 9.5\times 10^{-2})$
@@
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 5: Performance of the ESPD initialized with SPDs from [28] with $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
2.5 Implications for QKD
As a representative application, we analyze the implications of the ESPD framework applied to QKD.
In a QKD protocol characterized by a secure error threshold $e_{th}$ and a baseline error rate $e_{C}$ (excluding detector-induced errors), the minimal tolerable channel transmission rate $\gamma$, which defines the operational limit of a link, is conventionally approximated by[27]:
Since the ESPD paradigm can substantially suppress effective DCR $d$ while simultaneously enhancing the DE $\eta$, it enables a dramatic reduction in $\gamma$. This suppression of the minimal transmission threshold directly translates into an extended communication distance and increased robustness against channel loss [^3] The channel transmission rate $t$ is conventionally calculated by $t=10^{-\frac{\alpha l}{10}}$, where $\alpha$ is a fixed parameter mainly determined by channel characters and $l$ is the communication distance.. Furthermore, the enhancement of $\eta$ not only lowers the threshold $\gamma$ but also scales the achievable secret key rate at any fixed channel transmission, significantly improving the system throughput.
Compared to existing alternative strategies, such as the empty-signal-detection paradigm [27], the ESPD framework offers a distinct functional advantage. While the previous method also reduces the channel transmission threshold, it introduces additional sifting consumption that inherently degrades the efficiency. In contrast, the ESPD framework achieves performance gains by directly improving DE and reducing DCR, without the need for auxiliary sifting. Consequently, the ESPD paradigm provides a more efficient approach for improving QKD performance under practical constraints.
3 Discussion
In this section, we discuss the technological feasibility and practical robustness of the ESPD framework, as well as its limitations and directions for future work.
3.1 Effectiveness when the Initial SPD is Outside the Convergent Range
As evidenced by the Monte Carlo analysis, the steady-state performance of the ESPD is largely independent of the specific characteristics of the initial SPD, provided that it lies within a convergence basin. However, although the convergence interval is relaxed, some traditional SPDs might still fall outside this range. For example, the InGaAs/InP SPD designed in 2017 with a relatively lower DE ($\approx 27.5\%$) and lower DCR ($\approx 10^{-6}$)[29] is outside the range in which the ESPD dynamics converge to a high-performance fixed point when a uniform configuration of $(n_{s},k_{s})=(6,3)$ or $(n_{s},k_{s})=(8,4)$ is shared by each level.
Nevertheless, such SPDs can still be significantly enhanced by the ESPD framework through an appropriate selection of the first-level configuration parameters $(n_{1},k_{1})$. This strategic parameter choice allows states of the dynamical system described by Eq. (8) to be driven into the basin of attraction at the first step. For instance, selecting $(n_{1},k_{1})=(6,1)$ and $(n_{1},k_{1})=(8,1)$ results in $(\eta_{1},d_{1})=(85.5\%,3.4\times 10^{-3})$ and $(\eta_{1},d_{1})=(90.0\%,4.5\times 10^{-3})$, respectively, starting from the aforementioned SPD from [29], which successfully falls within the convergence basin for subsequent levels. Indeed, while a trade-off exists between DE and DCR, DCR of the ESPD can be drastically reduced by increasing $k$, and DE can be significantly enhanced by selecting a small $k$. Therefore, by employing one or two levels of the ESPD paradigm with suitable configuration parameters as a buffer, the detector’s performance can be adjusted into a desirable operational window, as long as the initial SPD is not prohibitively low.
Numerical simulations for an initial SPD with $(\eta_{0},d_{0})=(27.5\%,1.0\times 10^{-6})$ are provided in Table 3 and visualized in Fig.6, suggesting that SPDs with either low DE and low DCR [29], or moderate DE and high DCR [28], can serve as viable initializations.
Table 3: Performance of ESPD initialized with an SPD from [29], with initial parameters $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
@@ 3.1 Effectiveness when the Initial SPD is Outside the Convergent Range
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 6: Performance of ESPD initialized with an SPD from [29] with initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$, respectively.
3.2 Technological Feasibility and Robustness under Degraded Parameters
The primary technological advantage of the ESPD paradigm lies in its exclusive reliance on accessible, room-temperature components, effectively circumventing the stringent cryogenic requirements inherent to superconducting SPDs. The technological feasibility of components is discussed as follows.
3.2.1 State Preparation and Measurement
The architectural flexibility of the ESPD framework allows for various DOFs to be employed, enabling the selection of an easily manipulated DOF based on specific platforms. In practical implementations, this facilitates high-fidelity SPAM, ensuring that the error parameter $Q$ can be maintained at a low level. Specifically, $Q$ can be suppressed to the $10^{-4}$ level using standard optical setups and commercial platforms [43, 42], guaranteeing $Q\ll 1$.
@@ 3.2.2 Controlled Gates
High transmission efficiencies, for example, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [42], are achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, remains invariant under an increasing number of cascaded modules or ESPD levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
Table 4: Performance of an ESPD initialized with an SPD from [28], with $p$ varied from 0.80 to 0.96. Simulations are performed for $(n,k)=(5,2)$ at each level $L$, and all other settings remain consistent with those described in Section 2.4.
@@
Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$ with degraded pp. All other settings remain consistent with those described in Section 2.4.
@@
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD paradigm remains valid for such probabilistic operations (with slight modifications in analysis), the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.7 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss accumulated across approximately 10 controlled modules with 0.1 dB attenuation on each.). This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 42].
@@
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths. Contemporary room-temperature implementations for C-NOT demonstrate high performance, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 42]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryogenic components[38, 39, 40, 41]. However, it is also worth further testing the paradigm’s resilience under degraded parameters. Here, we also provide simulation results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), whose visual representation is shown in Fig.8 and 9. All other settings remain consistent with those described in Section 2.4. The numerical results indicate the robustness of the scheme: Comparison of Table 5 with Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded, such that $P=0.40$. This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with and near-term room-temperature quantum fabrication processes.
@@
Table 5: Performance evolution of the ESPD initialized with an SPD from [28], and moderate controlled gate fidelity, $P=0.80$.
@@
(a) DE $\eta_{s}$ vs. level $s$.
(b) DCR $d_{s}$ vs. level $s$.
Figure 8: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
Table 6: Performance evolution of the ESPD initialized with an SPD from [28], under a low controlled gate fidelity, $P=0.40$.
@@
(a) DE $\eta_{s}$ vs. level $s$.
(b) DCR $d_{s}$ vs. level $s$.
Figure 9: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
@@ 3.2.3 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single-photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, which is a global parameter regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more detection opportunities for the required state $|1\rangle$ [^4] Since the architecture incorporates a filter block, any additional $|0\rangle$ components will be filtered out. Therefore, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
3.3 Limitations and Future Work
@@ 3.3.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the Monte Carlo analysis and numerical simulations presented in Section 2.4 are sufficient for convergence demonstration and physical implementation guidance, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions and calculating the precise steady-state point from a more theoretical perspective. Given the inherent complexity of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
3.3.2 Resource Overhead vs. Optional Simplicity
@@
While the parallelized nature of the ESPD detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that increases exponentially with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although significant performance gains are typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling increases the architectural footprint and component count compared to a standalone SPD. For instance, a three-level ESPD scheme employing five controlled modules per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.2), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
@@
However, this complexity is compatible with modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], confirming that the integration density required for a multi-level ESPD is fully compatible with standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
@@
In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation. The SNSPD mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K). The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration. By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost. This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.